# A Survey for Foundation Models in Autonomous Driving

### v1

@@ 2.2.2 Prediction
Prediction forecasts traffic participants’ future trajectories, intents, and possible interactions with the ego vehicle. The common deep learning-based models are based on rasterized or vector images of the traffic scene, which encode spatial information. However, it is still challenging to accurately predict highly interactive scenes, which requires reasoning and semantic information, for example, right-of-ways, vehicles’ turning signals, and pedestrians’ gestures. The text representation of the scene can provide more semantic information, and better leverage LLM’s reasoning capability and common knowledge in the pre-training dataset. There is still not much research applying LLM to trajectory prediction. Keysan et al. (2023) did an early exploration of LLM’s power to make trajectory predictions. They convert the scene representation into text prompts, and use BERT model to generate the text encoding, which is finally fused with image encoding to decode trajectory prediction. Their evaluation shows significant improvement compared with baselines only using image encoding or text encoding.

### v3

@@ 2.2.2 Prediction
Prediction forecasts traffic participants’ future trajectories, intents, and possible interactions with the ego vehicle. The common deep learning-based models are based on rasterized or vector images of the traffic scene, which encode spatial information. However, it is still challenging to accurately predict highly interactive scenes, which requires reasoning and semantic information, for example, right-of-ways, vehicles’ turning signals, and pedestrians’ gestures. The text representation of the scene can provide more semantic context and better leverage LLM’s reasoning capability and common knowledge in the pre-training dataset.
Keysan et al. (2023) did an early exploration of LLM’s power to make trajectory predictions. They convert the scene representation into text prompts, and use BERTDevlin et al. (2018) model to generate the text encoding, which is finally fused with image encoding to decode trajectory prediction. Their evaluation shows significant improvement compared with baselines only using image encoding or text encoding.
LG-TrajChib and Singh (2024) leverages pre-trained LLM to generate motion cues for pedestrian trajectory prediction. Given the prompt and pedestrian coordinates, LLM is able to identify pedestrian motion patterns, such as linear motion, curved motion, and standing still. Combined with the Transformer-based motion encoder and social decoder, their method achieved the best results in multiple data sets compared with other state-of-the-art methods.
iMotion-LLMFelemban et al. (2024) proposes a novel motion prediction method leveraging instructions and captions in text. Given the text instruction inputs and scene encoding projected into text space, LLM decodes the output captions describing the driving decisions and latent tokens to generate multi-modal trajectories. Their model significantly outperforms the baseline model without using LLM. Additionally, they propose a feasibility classification task, which aligns predicted trajectories with feasible instructions and reject infeasible ones to improve safety.